The senior class at Richmont High School is selling t-shirts to raise money for its prom. The equation R(p) = - 25p2 + 600p describes the revenue, R, in dollar s, as a function of the price, p, in dollars, of a t-shirt . That is, the revenue is a function of price.
a) Determine the revenue if the group sells each shirt for $10.
b) Determine the revenue if the group sells each shirt for $15.
c) If the senior class hopes to have a revenue of
$3600, how much should it charge for each t-shirt?


Answer by Glaviolette(140)

a. R(p) = -25(10)^2 + 600(10) = $3500
b. R(p) = -25(15)^2 + 600(15) = $3375
c. -25p^2 + 600p = 3600
-25p^2 + 600p - 3600 = 0
-25 (p^2 - 24p + 144) = 0
-25 (p - 12)(p - 12) = 0
p-12 = 0
p = 12
$12